OpenStudy (anonymous):

Please check my answer Triangle ABC has been dilated to form triangle AꞌBꞌCꞌ.What is the scale factor?

OpenStudy (anonymous):

i think it is 2/5 once simplified

OpenStudy (anonymous):

@igreen

OpenStudy (anonymous):

@radar

OpenStudy (anonymous):

@Kainui

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

simply take the ratio of sides of image and preimage : \[\large \dfrac{7.5}{3}\]

OpenStudy (anonymous):

yes , I did and simplified I got 2/5

OpenStudy (anonymous):

I did the 7.5/3...3/7.5...becomes 2/5

ganeshie8 (ganeshie8):

you did it in reverse

ganeshie8 (ganeshie8):

7.5/3 should be 5/2

ganeshie8 (ganeshie8):

\[\large \dfrac{7.5}{3} = \dfrac{75}{30} = \dfrac{5}{2}\]

OpenStudy (anonymous):

I did it like this one... Pre-image ABCD was dilated to produce image AꞌBꞌCꞌDꞌ. What is the scale factor of the similar figures? Enter the answer in simplest form. Two quadrilaterals labeled A B C D and A-prime, B-prime, C-prime, D-prime. Quadrilateral A B C D is larger than the other quadrilateral, which is to the right of quadrilateral A B C D. Side B C is labeled twelve, and side B-prime, C-prime is labeled nine.

OpenStudy (anonymous):

it was explained to me that I needed to reverse the numerator and the denominator .... now confused :{

ganeshie8 (ganeshie8):

just a convention always put the "image" side on top

OpenStudy (anonymous):

ok... so i need to go back and fix the other as well then or it will be incorrect?.

ganeshie8 (ganeshie8):

whats your answer for the other one with quadrilateral ?

OpenStudy (anonymous):

simplified 3/4.... but that was doing the switching thing bc I originally had 4/3

ganeshie8 (ganeshie8):

3/4 is correct.

OpenStudy (anonymous):

ok, but what is the reason I flip the one and not the other? I just want to make sure I understand it

ganeshie8 (ganeshie8):

no wait, in the first question, the bigger triangle is image, yes ?

OpenStudy (anonymous):

let me reread it

OpenStudy (anonymous):

it reads as if the ABC is the original and the other is the dilation

OpenStudy (anonymous):

if it is image on top the the first would be 12/9 simplified to 4/3

ganeshie8 (ganeshie8):